Larry Mitchell invested part of his $21,000 advance at 8% annual simple interest and the rest at 7% annual simple interest. If his total yearly interest from both accounts was $1,580​, find the amount invested at each rate.
Answer by addingup(3677)

Some at 0.08
Some at 0.07
Total 21,000
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Let's call the investment at 8% x
What's left of the 21,000 is at 7%
So:
0.08 = x
0.07 = 21,000-x
0.08x+0.07(21,000-x) = 1,580
0.08x+1,470-0.07x = 1,580 Subtract x on left and subtract 1,470 on both sides:
0.01x = 110 Divide both sides by 0.01
x = 11,000 is at 8%. And:
21,000-11,000 = 10,000 at 7%
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
0.08*11,000 = 880
0.07*10,000 = 700
___________-------
Total... = 1,580 We have the correct answer